OpenStudy (anonymous):

f(x)=x^3=x^2 g(x)= x^3/2+5 Determine (F o G)(1)

OpenStudy (anonymous):

f(x)= x^3 + x^2

OpenStudy (anonymous):

(F o G) who is "o"?

OpenStudy (anonymous):

f(g(1))

OpenStudy (mertsj):

I am confused by f(x)=x^3=x^2 Is that what it is supposed to say?

hero (hero):

I think you posted f(x) incorrectly @jfry

OpenStudy (anonymous):

i did, it f(x)= x^3 + x^2

OpenStudy (mertsj):

Now who lectured me about exactly what he is doing?????

hero (hero):

I hope you're not referring to me. I didn't lecture you about anything

OpenStudy (mertsj):

\[f(g(x))=(x ^{\frac{3}{2}}+5)^{3}+(x ^{\frac{3}{2}}+5)^{2}\]

OpenStudy (anonymous):

Thank you

OpenStudy (mertsj):

However, if all you want is the f(g(1)), the simplest approach would be to find the g(1) and put that result into f

OpenStudy (mertsj):

g(1)=6 f(6)=216+36=252

OpenStudy (anonymous):

Im not sure im getting this

OpenStudy (mertsj):

The f(g(x)) means replace every x in the function f with the function g

OpenStudy (anonymous):

i got it now, i was having trouble with the x^3/2. i thought it would be 1/2, not 1.

OpenStudy (mertsj):

Sometimes it helps to make a diagram of the function f prepared to receive the function g